OpenStudy (anonymous):

The graph of a function is shown here. A coordinate plane shows five ordered pairs being plotted. From the origin, the first dot is zero units right and seven units up. The second dot is two units right and six units up. The third dot is four units right and five units up. The fourth dot is eight units right and three units up. The fifth dot is six units right and four units up. Which ordered pair, when added to the graph, will keep the graph a function? (4, 4) (6,1) (10, 2) (2, 9)

OpenStudy (anonymous):



OpenStudy (anonymous):

are all of those from the origin? or ontop of eachother

OpenStudy (anonymous):

find the equation of the line then plug in the points

OpenStudy (anonymous):

from the origin

OpenStudy (anonymous):

y=mx+b y intercept is 7 so b =7 slope=m= (y2-y1)/(x2-x1) use the points (0,7) and (2,6)

OpenStudy (anonymous):

so m=-1/2

OpenStudy (anonymous):

and y=-1/2x+7 plug in the possible values into this equation to see if they are valid

OpenStudy (anonymous):

example (4,4) y=-1/2x+7 (4)=-1/2(4)+7 4=-2+7 4=5 this doesn't make sense so (4,4) is not the point we want contine in this fashion until you find a point that works

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so (6, 1)?